# Multiple exponential sums and their applications to quadratic congruences

### v3

@@ Abstract.
In this paper, we develop a method of evaluating general exponential sums with rational amplitude functions for multiple variables which complements works by T. Cochrane and Z. Zheng on the single variable case. As an application, for $n\geq 3$, a fixed natural number, we obtain an asymptotic formula for the (weighted) number of solutions of general quadratic congruences modulo $p^{m}$ in small boxes, thus establishing an equidistribution result for these solutions.
@@ 1. Introduction and main results
In this paper, we develop a method for evaluating and estimating general exponential sums with rational amplitude functions for multiple variables. A general pure exponential sum (i.e., without twist by multiplicative characters) in $n$ variables is of type
@@ 1.1. Notations
    We assume that $p$ is an odd prime, $m\geq 2$ is an integer and $f$ is a rational function of the form $f_{1}/f_{2}$, where $f_{1},f_{2}\in\mathbb{Z}[x_{1},...,x_{n}]$ such that the reductions of $f_{1}$ and $f_{2}$ modulo $p$ are coprime.
@@
    The Hessian of a function $f$, denoted by Hesse$(f)$, is defined as
    $\text{Hesse}(f)=\begin{bmatrix}\frac{\partial^{2}f}{\partial x_{1}^{2}}&\frac{\partial^{2}f^{2}}{\partial x_{1}\partial x_{2}}&\cdots&\frac{\partial^{2}f}{\partial x_{1}\partial x_{n}}\\ \frac{\partial^{2}f}{\partial x_{2}x_{1}}&\frac{\partial^{2}f^{2}}{\partial x_{2}^{2}}&\cdots&\frac{\partial^{2}f}{\partial x_{2}\partial x_{n}}\\ \vdots&\vdots&\cdots&\vdots\\ \frac{\partial^{2}f}{\partial x_{n}\partial x_{1}}&\frac{\partial^{2}f^{2}}{\partial x_{n}\partial x_{2}}&\cdots&\frac{\partial^{2}f}{\partial x_{n}^{2}}\end{bmatrix}.$
@@
    $\text{ord}_{p}(f)=\text{ord}_{p}(f_{1})-\text{ord}_{p}(f_{2}),$
@@
    $\text{ord}_{p}(f_{i})=\max\{k\in\mathbb{Z}_{\geq 0}|~{}p^{k}~{}\text{divides all coefficients of}~{}f_{i}\}.$
@@
    If $r=\min_{1\leq i\leq n}\{\text{ord}_{p}(\frac{\partial f}{\partial x_{i}})\}$, we denote by $\mathcal{C}_{f}$ the set of critical points defined as
@@
    We denote by $G_{q}$ the quadratic Gauss sum
@@
    For any $n$-tuple $\bf{\alpha}$, we define
@@
Let $p$ be an odd prime and $f$ be a non-constant rational function in $n$ variables defined over $\mathbb{Z}$. Then we have
(1.1) $\begin{split}S_{\bf{\alpha}}(f,p^{m})=\begin{cases}p^{n(m+r)/2}\displaystyle\sum_{\begin{subarray}{c}{\bf{y}}\in(\mathbb{Z}p/p^{(m-r)/2}\mathbb{Z}_{p})^{n}\\ {\bf{y}}\equiv{\bf{\alpha}}\bmod p\\ p^{-r}\nabla f({\bf{y}})\equiv 0\bmod p^{(m-r)/2}\end{subarray}}e_{p^{m}}(f({\bf{y}})),~{}&m-r~{}\text{is even};\\ p^{(m+r-1)n/2}\displaystyle\sum_{\begin{subarray}{c}{\bf{y}}\in(\mathbb{Z}p/p^{(m-r+1)/2}\mathbb{Z}_{p})^{n}\\ {\bf{y}}\equiv{\bf{\alpha}}\bmod p\\ p^{-r}\nabla f({\bf{y}})\equiv 0\bmod p^{(m-r-1)/2}\end{subarray}}e_{p^{m}}(f({\bf{y}})),~{}&m-r~{}\text{is odd}.\end{cases}\end{split}$
@@
Let $p$ be an odd prime and $f$ be a non-constant rational function in $n$ variables defined over $\mathbb{Z}$. Then we have the following.
@@
    $S_{\bf{\alpha}}(f,p^{m})=\begin{cases}p^{n(m+r)/2}e_{p^{m}}(f({\bf{{\bf{\alpha}^{*}}}})),~{}&m-r~{}\text{\rm is even};\\ \left(\frac{\text{\rm det}(p^{-r}\text{\rm Hesse}f({\bf{\alpha}^{*}}))}{p}\right)p^{(m+r-1)n/2}e_{p^{m}}(f({\bf{\alpha}^{*}}))G_{p}(1)^{n},~{}&m-r~{}\text{\rm is odd},\end{cases}$
    where $\bf{\alpha}^{*}$ is the unique lifting of a solution $\alpha$ of the congruence $p^{-r}\nabla f({\bf x})={\bf{0}}\bmod p$ to a solution of the congruence $p^{-r}\nabla f({\bf{x}})\equiv 0\bmod p^{[\frac{m-r+1}{2}]}$.
Remark 1.
In the situation of Theorem 1, if $y\equiv\alpha^{*}\bmod p^{{[\frac{m-r+1}{2}]}},$ then $e_{p^{m}}(f({\bf{y}}))=e_{p^{m}}(f(\alpha^{*})).$ In the subsequent part of this article, by $\alpha^{*}$ we mean the unique lifting of $\alpha$ to a solution of the congruence $p^{-r}\nabla f({\bf{x}})\equiv 0\bmod p^{{m-r}}$.
@@
Based on Theorem 1, which will be enough for our application, we will derive an asymptotic formula for the number of solutions of special quadratic congruences in several variables modulo prime powers in small boxes. In the following, we give some motivation for this problem.
@@
[…]
@@
Here we assume that $A=(a_{ij})_{1\leq i,j\leq n}$ is a symmetric matrix such that $(\det(A),p)=1$, i.e. $Q$ is non-singular modulo $p$. We count weighted solutions in boxes $I_{1}\times\cdots\times I_{n}$, where $I_{1},...,I_{n}$ are intervals of length $2N$, which we aim to take as small as possible compared to the modulus. The main novelties in this paper, as compared to the previous works [2] and [13], are two-fold. Firstly, we now use evaluations of multivariable exponential sums, as provided in Proposition 1 and Theorem 1. Secondly, we directly use the estimate for exponential sums to get the number of solutions of the general quadratic form. This simplifies our evaluations a lot, and the method becomes more elegant.
@@
$Q(x_{1},x_{2},..x_{n})=0$
since in this case the squares of the variables are much smaller than the modulus $q$. Here the number of solutions satisfies an asymptotic of the form $\sim c_{p}N^{n-2}\log N$ for a certain constant $c_{p}$ depending on $p$. In particular, in the case $n=2$, we count Pythagorean triples. We point out that the case of Pythagorean triples modulo prime powers was considered in the unpublished work [3]. In [13] the third-named author studied the general quadratic form for $n=3$ with some restrictions to the coefficients of the form. In this paper, we study arbitrary non-singular quadratic from for $n$ variables without any such additional restrictions. This we establish in the next theorem.
@@
Let $\varepsilon>0$ and $n(\geq 3)\in\mathbb{N}$ be fixed, $\Phi:\mathbb{R}\rightarrow\mathbb{R}_{\geq 0}$ be a Schwartz class function and $p$ be an odd prime. Let $(x_{0,1},x_{0,2},...,x_{0,n})$ be a fixed point in $\mathbb{Z}^{n}$. Then as $m\rightarrow\infty$, we have the asymptotic formula
@@
$\mathcal{A}$ =$=\left\{(x_{1},...,x_{n})\in\mathbb{Z}^{n}|~{}1\leq x_{1},...,x_{n}\leq p^{m},~{}(x_{1},...,x_{n})\not\equiv(0,...,0)\bmod p,\ \sum_{i,j=1}^{n}a_{ij}x_{i}x_{j}\equiv 0\bmod p^{m}\right\}.$
@@
Note that $b_{p}$ is independent of $m$, which can be easily seen using Proposition 1 of [29]. Moreover, looking at the case $m=1$, we have $b_{p}>0$ as a consequence of [24, subsection 1.7, Proposition 4].
Our proof of Theorem 2 above follows the main lines in [2] and [13], but we proceed without using parametrization. Instead we directly apply Poisson summation and evaluation of exponential sums to prove our result. However, it contains significant novelties which simplify the treatment, as described above.
Acknowledgements. The authors would like to thank Ramakrishna Mission Vivekananda Educational and Research Institute for providing excellent working conditions. A.H. would like to thank Professor S.D. Adhikari for his suggestion to consider quadratic forms of $n$ variables. During the preparation of this article, N.B. was supported by the National Board of Higher Mathematics post-doctoral fellowship (No.: 0204/3/2021/R&D-II/7363) and A.H. would like to thank CSIR, Government of India for financial support in the form of a Senior Research Fellowship (No.: 09/934(0016)/2019-EMR-I).
@@ 2. Preliminaries
This is carried out in the following proposition.
Proposition 3.
For any $\varepsilon>0$ and a non-singular quadratic form $Q(x_{1},...,x_{n})=\sum_{i,j=1}^{n}b_{ij}x_{i}x_{j}$, where $b_{ij}\in\mathbb{Z}$, we have
$\sharp\{(x_{1},x_{2},...,x_{n})\in\mathbb{Z}^{n}:Q(x_{1},...,x_{n})=0,\ |x_{i}|\leq N\}=O_{\varepsilon}(N^{n-2+\varepsilon}),$
as $N\rightarrow\infty$.
@@ 3. Proof of Proposition 1
Let $m-r$ be even. Writing ${\bf{x}}=p^{(m-r)/2}{\bf{z}}+{\bf{y}}$ with ${\bf z}\in\left(\mathbb{Z}_{p}/p^{(m+r)/2}\mathbb{Z}_{p}\right)^{n}$ and ${\bf y}\in\left(\mathbb{Z}_{p}/p^{(m-r)/2}\mathbb{Z}_{p}\right)^{n}$, we obtain the following chain of equations.
$S_{\bf{\alpha}}(f,p^{m})$ $=\sum_{\begin{subarray}{c}{\bf{x}}\in(\mathbb{Z}p/p^{m}\mathbb{Z}_{p})^{n}\\ {\bf{x}}\equiv{\bf{\alpha}}\bmod p\end{subarray}}e_{p^{m}}(f({\bf{x}}))$
$=\sum_{{\bf{z}}\in(\mathbb{Z}p/p^{(m+r)/2}\mathbb{Z}_{p})^{n}}\sum_{\begin{subarray}{c}{\bf{y}}\in(\mathbb{Z}p/p^{(m-r)/2}\mathbb{Z}_{p})^{n}\\ {\bf{y}}\equiv{\bf{\alpha}}\bmod p\end{subarray}}e_{p^{m}}(f({p^{(m-r)/2}\bf{z}}+{\bf y}))$
$=\sum_{\begin{subarray}{c}{\bf{y}}\in(\mathbb{Z}p/p^{(m-r)/2}\mathbb{Z}_{p})^{n}\\ {\bf{y}}\equiv{\bf{\alpha}}\bmod p\end{subarray}}\sum_{{\bf{z}}\in(\mathbb{Z}p/p^{(m+r)/2}\mathbb{Z}_{p})^{n}}e_{p^{m}}(f({\bf{y}})+p^{(m-r)/2}{\bf{z}}\cdot\nabla f({\bf{y}})+p^{m-r}2^{-1}{\bf{z}}^{t}\text{Hesse}~{}f({\bf{y}}){\bf{z}})$
$=\sum_{\begin{subarray}{c}{\bf{y}}\in(\mathbb{Z}p/p^{(m-r)/2}\mathbb{Z}_{p})^{n}\\ {\bf{y}}\equiv{\bf{\alpha}}\bmod p\end{subarray}}e_{p^{m}}(f({\bf{y}}))\sum_{{\bf{z}}\in(\mathbb{Z}p/p^{(m+r)/2}\mathbb{Z}_{p})^{n}}e_{p^{m}}(p^{(m-r)/2}{\bf{z}}\cdot\nabla f({\bf{y}}))$
$=p^{nr}\sum_{\begin{subarray}{c}{\bf{y}}\in(\mathbb{Z}p/p^{(m-r)/2}\mathbb{Z}_{p})^{n}\\ {\bf{y}}\equiv{\bf{\alpha}}\bmod p\end{subarray}}e_{p^{m}}(f({\bf{y}}))\sum_{{\bf{z}}\in(\mathbb{Z}p/p^{(m-r)/2}\mathbb{Z}_{p})^{n}}e_{p^{(m-r)/2}}(p^{-r}{\bf{z}}\cdot\nabla f({\bf{y}}))$
$=p^{n(m+r)/2}\sum_{\begin{subarray}{c}{\bf{y}}\in(\mathbb{Z}p/p^{(m-r)/2}\mathbb{Z}_{p})^{n}\\ {\bf{y}}\equiv{\bf{\alpha}}\bmod p\\ p^{-r}\nabla f({\bf{y}})\equiv 0\bmod p^{(m-r)/2}\end{subarray}}e_{p^{m}}(f({\bf{y}})).$
Note that in the Taylor development in the second line, the $l$-th order term of the Taylor polynomial vanishes modulo $p^{m}$ for $l\geq 3$. Indeed, for $l\geq 3$ and $r\leq m-2$, the order of $p$ is equal to ${{\frac{(m-r)l}{2}}-\sum_{k=1}^{\infty}[\log_{p^{k}}l]}-(m-r)$, and

### v5

@@ Abstract.
In this paper, we develop a method of evaluating general exponential sums with rational amplitude functions for multiple variables which complements work by T. Cochrane and Z. Zheng on the single variable case. As an application, for $n\geq 3$, a fixed natural number, we obtain an asymptotic formula for the (weighted) number of solutions of general quadratic congruences modulo $p^{m}$ in small boxes, thus establishing an equidistribution result for these solutions.
@@ 1. Introduction and main results
In this paper, we develop a method for evaluating and estimating general exponential sums with rational amplitude functions for multiple variables. A general pure exponential sum (i.e., without twist by multiplicative characters) in $n$ variables is of the type
@@ 1.1. Notations
    We assume that $p$ is an odd prime, $m\geq 2$ is an integer and $f$ is a rational function of the form $f_{1}/f_{2}$, where $f_{1},f_{2}\in\mathbb{Z}[x_{1},...,x_{n}]$ are such that the reductions of $f_{1}$ and $f_{2}$ modulo $p$ are coprime.
@@
    The Hessian of a function $f$, denoted by Hesse(f)(f), is defined as
    $\text{Hesse}(f)=\begin{bmatrix}\frac{\partial^{2}f}{\partial x_{1}^{2}}&\frac{\partial^{2}f}{\partial x_{1}\partial x_{2}}&\cdots&\frac{\partial^{2}f}{\partial x_{1}\partial x_{n}}\\ \frac{\partial^{2}f}{\partial x_{2}x_{1}}&\frac{\partial^{2}f}{\partial x_{2}^{2}}&\cdots&\frac{\partial^{2}f}{\partial x_{2}\partial x_{n}}\\ \vdots&\vdots&\cdots&\vdots\\ \frac{\partial^{2}f}{\partial x_{n}\partial x_{1}}&\frac{\partial^{2}f}{\partial x_{n}\partial x_{2}}&\cdots&\frac{\partial^{2}f}{\partial x_{n}^{2}}\end{bmatrix}$.
@@
    $\text{\rm ord}_{p}(f)=\text{\rm ord}_{p}(f_{1})-\text{\rm ord}_{p}(f_{2}),$
@@
    $\text{\rm ord}_{p}(f_{i})=\max\{k\in\mathbb{Z}_{\geq 0}|~p^{k}~\text{divides all coefficients of}~f_{i}\}.$
@@
    If $r=\min_{1\leq i\leq n}\{\text{\rm ord}_{p}(\frac{\partial f}{\partial x_{i}})\}$, we denote by $\mathcal{C}_{f}$ the set of critical points defined as
@@
    We denote by $G_{q}(n)$ the quadratic Gauss sum
@@
    For any $n$-tuple ${\alpha}\in\mathbb{Z}^{n}$, we define
@@
  • 10.
    Throughout the article $\overline{x}$ denotes the multiplicative inverse of $x$ to relevant modulus.
@@
Let $p$ be an odd prime and $f$ be a non-constant rational function in $n$ variables defined over $\mathbb{Z}$. Set $r:=\min_{1\leq i\leq n}\{\text{\rm ord}_{p}(\frac{\partial f}{\partial x_{i}})\}$. Then we have
(1.1) $\begin{split}S_{\bf{\alpha}}(f,p^{m})=\begin{cases}p^{n(m+r)/2}\displaystyle\sum_{\begin{subarray}{c}{\bf{y}}\in(\mathbb{Z}_{p}/p^{(m-r)/2}\mathbb{Z}_{p})^{n}\\ {\bf{y}}\equiv{\bf{\alpha}}\bmod p\\ p^{-r}\nabla f({\bf{y}})\equiv 0\bmod p^{(m-r)/2}\end{subarray}}e_{p^{m}}(f({\bf{y}})),~&m-r~\text{is even};\\ p^{(m+r-1)n/2}\displaystyle\sum_{\begin{subarray}{c}{\bf{y}}\in(\mathbb{Z}_{p}/p^{(m-r+1)/2}\mathbb{Z}_{p})^{n}\\ {\bf{y}}\equiv{\bf{\alpha}}\bmod p\\ p^{-r}\nabla f({\bf{y}})\equiv 0\bmod p^{(m-r-1)/2}\end{subarray}}e_{p^{m}}(f({\bf{y}})),~&m-r~\text{is odd}.\end{cases}\end{split}$
@@
Let $p$ be an odd prime and $f$ be a non-constant rational function in $n$ variables defined over $\mathbb{Z}$. Let $m\in\mathbb{N}$, define $r$ as above and assume that $m-r\geq 2$. Then we have the following.
@@
    $S_{\bf{\alpha}}(f,p^{m})=\begin{cases}p^{n(m+r)/2}e_{p^{m}}(f({\bf{{\bf{\alpha}^{*}}}})),~&m-r~\text{\rm is even};\\ \left(\frac{\text{\rm det}(\overline{2}p^{-r}\text{\rm Hesse}(f)({\bf{\alpha}^{*}}))}{p}\right)p^{(m+r-1)n/2}e_{p^{m}}(f({\bf{\alpha}^{*}}))G_{p}(1)^{n},~&m-r~\text{\rm is odd},\end{cases}$
    where $\overline{2}$ is a multiplicative inverse of $2$ modulo $p^{m}$ and $\bf{\alpha}^{*}$ is the unique lifting of a solution $\alpha$ of the congruence $p^{-r}\nabla f({\bf x})={\bf{0}}\bmod p$ to a solution of the congruence $p^{-r}\nabla f({\bf{x}})\equiv 0\bmod p^{m}$.
@@
Since complete exponential sums with rational functions appear frequently in analytic number theory, it may be expected that Theorem 1 can be applied to many situations. In this article, we will employ it to derive an asymptotic formula for the number of solutions of special quadratic congruences in several variables modulo prime powers in small boxes. Here the relevant exponential sums are quadratic Gauss sums in several variables with certain restrictions on the summation variables. Closely related exponential sums were evaluated explicitly in [17, sections 10] for prime and prime square moduli. We infer our evaluations for arbitrary prime power moduli in a straight-forward way from the much more general Theorem 1 above. In the following, we give some motivation for the problem under consideration.
@@
modulo prime powers $q=p^{m}$ with an arbitrary sequence $\alpha=(\alpha_{1},...,\alpha_{n})$ of coefficients (i.e. $\alpha$ not depending on $m$), the second- and third-named authors [2] improved the said exponent $2/3$ to $11/18$ under the stronger condition $(x_{1}\cdots x_{n},p)=1$ on the solutions $(x_{1},...,x_{n})$ of (1.2) (see [2, Theorem 2]). A result by Cochrane [9] for general moduli $q$ implies that for any form $Q(x)$ with $\it fixed$} coefficients, there is a non-zero solution with $\max\{|x_{1}|,|x_{2}|,|x_{3}|\}=O(q^{1/2})$, where the $O$-constant may depend on the form. In the above-mentioned paper [2], the second- and third-named authors refined this result into an asymptotic formula for the case of fixed diagonal forms modulo prime powers (see [2, Theorem 1]). This was generalized by the third-named author in [13] to general quadratic forms $ax^{2}+bxy+cy^{2}+dxz+eyz+fz^{2}$ with fixed coefficients.
@@
Here we assume that $A=(a_{ij})_{1\leq i,j\leq n}$ is a symmetric matrix such that $(\det(A),p)=1$, i.e. $Q$ is non-singular modulo $p$. We count weighted solutions in boxes $I_{1}\times\cdots\times I_{n}$, where $I_{1},...,I_{n}$ are intervals of length $2N$, which we aim to take as small as possible compared to the modulus. The main novelties in this paper, as compared to the previous works [2] and [13], are two-fold. Firstly, we now use evaluations of multivariable exponential sums, as provided in Proposition 1 and Theorem 1. Secondly, we infer our results directly from estimates for these exponential sums, avoiding a parametrization of solutions of quadratic congruences which was used in [2] and [13]. This simplifies our evaluations a lot, and the method becomes more elegant.
@@
$Q(x_{1},x_{2},...,x_{n})=0$
since in this case the squares of the variables are much smaller than the modulus $q$. Here the number of solutions satisfies an asymptotic of the form $\sim cN^{n-2}(\log N)^{r_{n}}$, where $c$ is a product of local densities and $r_{n}=1$ if $n\leq 4$ and $r_{n}=0$ if $n>4$. In particular, in the case $n=2$, we count Pythagorean triples. We point out that the case of Pythagorean triples modulo prime powers was considered in the unpublished work [3]. In [13] the third-named author studied the general quadratic form for $n=3$ with some restrictions to the coefficients of the form. In this paper, we study arbitrary non-singular quadratic from for $n$ variables without any such additional restrictions. We establish the following.
@@
Assume that $A=(a_{ij})_{1\leq i,j\leq n}$ is a symmetric matrix with integer coefficients such that $(\det(A),p)=1$, i.e. $Q$ is non-singular modulo $p$. Let $\varepsilon>0$ and $n(\geq 3)\in\mathbb{N}$ be fixed, $\Phi:\mathbb{R}\rightarrow\mathbb{R}_{\geq 0}$ be a Schwartz class function and $p$ be an odd prime. Let $(x_{0,1},x_{0,2},...,x_{0,n})$ be a fixed point in $\mathbb{Z}^{n}$. Then as $m\rightarrow\infty$, we have the asymptotic formula
@@
$\mathcal{A}$ $=\left\{(x_{1},...,x_{n})\in\mathbb{Z}^{n}|~1\leq x_{1},...,x_{n}\leq p^{m},~(x_{1},...,x_{n})\not\equiv(0,...,0)\bmod p,\ \sum_{i,j=1}^{n}a_{ij}x_{i}x_{j}\equiv 0\bmod p^{m}\right\}.$
@@
Note that $b_{p}$ is independent of $m$. We have
$\sharp\mathcal{A}=p^{(n-1)(m-1)}\sharp\left\{(x_{1},...,x_{n})\in\mathbb{Z}^{n}|~1\leq x_{1},...,x_{n}\leq p,~(x_{1},...,x_{n})\not\equiv(0,...,0)\bmod p,\ \sum_{i,j=1}^{n}a_{ij}x_{i}x_{j}\equiv 0\bmod p\right\},$
which can be shown using a Hensel’s type argument. To elaborate more, assume that $(x_{1}+k_{1}p^{k},...,x_{n}+k_{n}p^{k})$ is a solution of $Q(x_{1},...,x_{n})=\sum_{i,j=1}^{n}a_{ij}x_{i}x_{j}$ modulo $p^{k+1}$, where $(x_{1},...,x_{k})$ is a solution modulo $p^{k}$ and $0\leq k_{i}\leq p-1$. Using Taylor expansion, $Q(x_{1}+k_{1}p^{k},...,x_{n}+k_{n}p^{k})=Q(x_{1},...,x_{n})+p^{k}(k_{1},...,k_{n})\cdot\nabla Q(x_{1},...,x_{n})\bmod p^{k+1}$. This gives
$(k_{1},...,k_{n})\cdot\nabla Q(x_{1},...,x_{n})\equiv-\frac{Q(x_{1},...,x_{n})}{p^{k}}\bmod p,$
which has $p^{n-1}$ number of solutions in $(k_{1},...,k_{n})$. This proves our claim by lifting up to modulo $p^{m}$. We have $b_{p}>0$ as a consequence of [24, subsection 1.7, Proposition 4].
Our proof of Theorem 2 above follows the main lines in [2] and [13], but we proceed without using parametrization. Instead we directly apply Poisson summation and evaluate the resulting exponential sums explicitly to prove our result.
Acknowledgements. The authors would like to thank the anonymous referees for their valuable comments. They would also like to thank the Ramakrishna Mission Vivekananda Educational and Research Institute for providing excellent working conditions. A.H. would like to thank Professor S.D. Adhikari for his suggestion to consider quadratic forms of $n$ variables. During the preparation of this article, N.B. was supported by the National Board of Higher Mathematics post-doctoral fellowship (No.: 0204/3/2021/R&D-II/7363) and A.H. would like to thank CSIR, Government of India for financial support in the form of a Senior Research Fellowship (No.: 09/934(0016)/2019-EMR-I).
Data availability statement. Data sharing is not applicable to this article as no new data were created or analyzed in this study.
Conflict of interest statement. The authors declare no conflicts of interest regarding this manuscript.
@@ 2. Preliminaries
This is carried out in the following proposition which is a consequence of a result in [17].
Proposition 3 (Heath-Brown).
Let $n\geq 3$ be an integer and $\varepsilon>0$. Then for any non-singular quadratic form $Q(x_{1},...,x_{n})=\sum_{i,j=1}^{n}b_{ij}x_{i}x_{j}$ with integer coefficients, we have
$\sharp\{(x_{1},x_{2},...,x_{n})\in\mathbb{Z}^{n}:Q(x_{1},...,x_{n})=0,\ |x_{i}|\leq N\}=O(N^{n-2+\varepsilon}),$
as $N\rightarrow\infty$, where the implied constant depends on $Q$ and $\varepsilon$.
@@ 3. Proof of Proposition 1
Let mrm-r be even. Writing ${\bf{x}}=p^{(m-r)/2}{\bf{z}}+{\bf{y}}$ with ${\bf z}\in\left(\mathbb{Z}_{p}/p^{(m+r)/2}\mathbb{Z}_{p}\right)^{n}$ and ${\bf y}\in\left(\mathbb{Z}_{p}/p^{(m-r)/2}\mathbb{Z}_{p}\right)^{n}$, we obtain the following chain of equations.
$S_{\bf{\alpha}}(f,p^{m})$ $=\sum_{\begin{subarray}{c}{\bf{x}}\in(\mathbb{Z}_{p}/p^{m}\mathbb{Z}_{p})^{n}\\ {\bf{x}}\equiv{\bf{\alpha}}\bmod p\end{subarray}}e_{p^{m}}(f({\bf{x}}))$
$=\sum_{{\bf{z}}\in(\mathbb{Z}_{p}/p^{(m+r)/2}\mathbb{Z}_{p})^{n}}\sum_{\begin{subarray}{c}{\bf{y}}\in(\mathbb{Z}_{p}/p^{(m-r)/2}\mathbb{Z}_{p})^{n}\\ {\bf{y}}\equiv{\bf{\alpha}}\bmod p\end{subarray}}e_{p^{m}}(f({p^{(m-r)/2}\bf{z}}+{\bf y}))$
$=\sum_{\begin{subarray}{c}{\bf{y}}\in(\mathbb{Z}_{p}/p^{(m-r)/2}\mathbb{Z}_{p})^{n}\\ {\bf{y}}\equiv{\bf{\alpha}}\bmod p\end{subarray}}\sum_{{\bf{z}}\in(\mathbb{Z}_{p}/p^{(m+r)/2}\mathbb{Z}_{p})^{n}}e_{p^{m}}(f({\bf{y}})+p^{(m-r)/2}{\bf{z}}\cdot\nabla f({\bf{y}})+\overline{2}p^{m-r}{\bf{z}}^{t}\text{Hesse}(f)({\bf{y}}){\bf{z}})$
$=\sum_{\begin{subarray}{c}{\bf{y}}\in(\mathbb{Z}_{p}/p^{(m-r)/2}\mathbb{Z}_{p})^{n}\\ {\bf{y}}\equiv{\bf{\alpha}}\bmod p\end{subarray}}e_{p^{m}}(f({\bf{y}}))\sum_{{\bf{z}}\in(\mathbb{Z}_{p}/p^{(m+r)/2}\mathbb{Z}_{p})^{n}}e_{p^{m}}(p^{(m-r)/2}{\bf{z}}\cdot\nabla f({\bf{y}}))$
$=p^{nr}\sum_{\begin{subarray}{c}{\bf{y}}\in(\mathbb{Z}_{p}/p^{(m-r)/2}\mathbb{Z}_{p})^{n}\\ {\bf{y}}\equiv{\bf{\alpha}}\bmod p\end{subarray}}e_{p^{m}}(f({\bf{y}}))\sum_{{\bf{z}}\in(\mathbb{Z}_{p}/p^{(m-r)/2}\mathbb{Z}_{p})^{n}}e_{p^{(m-r)/2}}(p^{-r}{\bf{z}}\cdot\nabla f({\bf{y}}))$
$=p^{n(m+r)/2}\sum_{\begin{subarray}{c}{\bf{y}}\in(\mathbb{Z}_{p}/p^{(m-r)/2}\mathbb{Z}_{p})^{n}\\ {\bf{y}}\equiv{\bf{\alpha}}\bmod p\\ p^{-r}\nabla f({\bf{y}})\equiv 0\bmod p^{(m-r)/2}\end{subarray}}e_{p^{m}}(f({\bf{y}})).$
Note that in the Taylor development in the third line, the $l$-th order term of the Taylor polynomial vanishes modulo $p^{m}$ for $l\geq 3$. Indeed, for $l\geq 3$ and $r\leq m-2$, the order of $p$ is equal to ${{\frac{(m-r)l}{2}}-\sum_{k=1}^{\infty}[\log_{p^{k}}l]}-(m-r)$, and